Previous

Calendar

April

2006

Next
Sunday Monday Tuesday Wednesday Thursday Friday Saturday

26

JD85

27

JD86

28

JD87

29

JD88

30

JD89

31

JD90
CU Holiday (Spring Break)

Site Visit (JH, MP, AT): Details

slow-response ozone instrument turned on (AT)

1

JD91

2

JD92
Change from MST to daylight savings time (MDT).

3

JD93

4

JD94

5

JD95
Site Visit (KR, MW): Details

160L LN2 Dewar is empty. Laser LN2 not filled.

soils sampled (MW).

6

JD96

7

JD97

8

JD98

9

JD99

10

JD100 (blue skies)
Site Visit (SpB, JH, JB): Details , Photos (30)

from 12:30-14:15 MST, Moved check-valve for LI-6262 down to pump, changed out copper heat exchanger.

4-flask swap at East/West Robots

11

JD101

12

JD102

13

JD103
(Full "Pink" Moon)

14

JD104
Site Visit (AT): Details

15

JD105

16

JD106
Easter

17

JD107

18

JD108

19

JD109 (clear/cool)
Site Visit (SpB, JH, MP, MW, M, NM): Details , Photos (26)

from 10:35-10:40 MST, swiped the Krypton Hyg.

from 11:10-12:00 MST, removed the loaner LI-6262 (sn IRG3-0224), and re-installed the Monson re-calibrated LI-6262 (sn IRG3-638)

at 12:54 MST, made slight adjustment to LI-6262 zero for co2/h2o.

from 13:06-13:23 MST, Changed profiler (li-6251) dessicant

from 13:30-13:40 MST, replaced N2 cylinder (new cyl= CC17728 at 1900 psi).

4-flask swap at East/West Robots

Snow spectrometer measurements (NM)

Prius Rental

20

JD110

21

JD111

22

JD112
Earth Day.

23

JD113

24

JD114 (snow)

25

JD115

26

JD116 (Cu/warm)
Site Visit (SpB, JH, SS): Details , Photos (34)

Transported three cylinders of N2 out to the tower.

No flasks swapped (but East Robot seems ok now)

Sean S. added more inlet levels to the TDL

Brought 3 empty cylinders back to campus.

Jeep Rental (drove up to C1)

27

JD117
Sean S. working on TGA

at 22:10 MST, quacker data system stopped.

28

JD118
Site Visit (SpB, ML, SS): Details , Photos (22)

Sean S. finished work on TGA

at around 1:00 MDT, russter2 dropped off network, power to the tower and trailer was down.

at around 12:00 MST, power to trailer restored (thanks to ML/MO/KC)

Tower Power Down

29

JD119
Tower Power Down

30

JD120
Tower Power Down

1

JD121
Site Visit (SpB, JL): Details , Photos (37)

10AM, met with John L from rainbow elec.

at 11:30 MST, Power Restored to Tower

Transformer at C1 has a short-circuit---bypassed C1 and tower transformers to restore power.

Brought 4 empty cylinders back to campus.

Jeep Rental

2

JD122
Site Visit (SpB): Details , Photos (21)

at around 3:00 MDT, russter2 dropped off network, the data system stopped at 1:42AM..

power restored to trailer, tower power is off (indefinitely).

3

JD123
Site Visit (SpB, JL): Details , Photos (14)

10AM, met with John L from rainbow elec.

Tracked down the short-circuit: it's in a j-box which is near where the conduit passes the trail (about 200' from CU tower). Poor weather and a lack of time stopped us from reconnecting everything.

Tower Power Down

4

JD124
Tower Power Down

5

JD125
Tower Power Down

6

JD126
Tower Power Down